Can someone help me solve this problem?
16 divided by (-2) * (-4)/2^3 + 4
Thanks,
Sher

Answer by fractalier(6550)

Looking at the order of operations PEMDAS, we get (and assuming it is written like this)
16 / [(-2) * (-4)/2^3 + 4] =
16 / [(-2) * (-4)/8 + 4] =
16 / [(-2) * (-1/2) + 4] =
16 / [1 + 4] =
16 / 5